Help is 4??b. A founder event, using the founder event as the reference generation and calculating the effective sizes 10 generations la5. Consider the pedigree below. What is the pedigree inbreeding coefficient F of the female shown by the solid circle? (5 pts
Answer #1

4. i) p = 0.1, f = 0.5 L q = 1-p = 1-0-1=0.9 SOT f CAA) = p² (1-4) + pf = 0.01 (1-0.5) + (0.18 0.5) = 0.005 +0.05 = 0.055 t (f(aa) = q l-ft of all = 0.04 C 1-0. 14/0.280.1) = 0.04x0.9 +0.02 = 0.036 +0.02 = 0.056 ili) p = 0.3, f = 0.05 te 9 = 1-6 - 1-= 0.09 C 1-0.05) + (0.3X -oios) = 0.09 1.05 -0.015 = 0.0945 -0.015 .= 0.0795 Alta) - 2pg (1-1) = 2x0.3 x 8.7 X (1--0.05). = 0

Thanks...